OpenStudy (anonymous):

The distance between two cities is ninety miles, and a woman drives from one city to the other at a rate of 45 mph. At what rate must she return if the total travel time is three hours and forty minutes?

OpenStudy (anonymous):

the distance between the two cities is 90 miles time = distance /speed 90 so, while driving at the speed of 45 miles/h, time taken to cover 90 miles = ----- = 2h 45 this means that the return trip has to be completed in 1 h 40 min = 100/60 h and distance is still 90 miles speed = distance/time 90 60 speed = -------- = 90 x ----- = 5400/100 = 54 mph 100 100 ---- 60 so the speed of the return journey has to be 54 mph